OpenStudy (anonymous):

Graph and find intersecting points for -9x+13y=3 and 2x-3y=-1. thanks

OpenStudy (anonymous):

What are the specifics? Are you allowed to use graphing utilities? What level of math is this? The route you take will depend on what you're familiar with. There's a few ways to do this.

OpenStudy (anonymous):

you solve the equation by putting it into slope intercept form. then you graph both equations and and then you get the point intersection its algebra 2

OpenStudy (anonymous):

Right, well depending on your level of math there are several other methods. Why I asked, is because this is such a nasty equation, none of the intercepts or slopes are very nice at all. Really ugly fractions. So first get each one into slope intercept form (y=mx+b)

OpenStudy (anonymous):

Looking at the first equation, can you get it into y=mx+b form?

OpenStudy (anonymous):

yeah i got both of them into slope intercept form i just need to graph them. but thats where im having trouble with. i have the correct answer which is (4.3) where the intersection is. but when i graph them its soomething way off. -_-

OpenStudy (anonymous):

y=9/13x+3/13 y=2/3x+3

OpenStudy (anonymous):

Yea, this is a really really poor question, I must say. The intercepts are all fractions, not nice ones at that, and all of the slopes are really nasty fractions. I don't think you can get an accurate answer on graph paper, I'm sorry :( You would either have to solve empirically or use a graphing utility such as a graphing calculator for an accurate answer.

OpenStudy (anonymous):

okay thanks.

OpenStudy (anonymous):

But, you should be able to get a "guesstimate" with graphing.

OpenStudy (anonymous):

i guess thats what ill do

OpenStudy (anonymous):

I'm afraid that is the best you CAN do with that type of question :/